Calculate the value of a preferred stock with a fixed annual dividend of $2.45, assuming a discount rate of 9.5%. Solve the problem two different ways: first by using the algebraic formula for a constant dividend preferred stock, then by using the built-in Excel function PV. hint: Use the Preferred Stock example in the posted DDM Excel Examples file as a guide. Feel free to copy the worksheet and make the minor necessary changes to answer this question.
2. Calculate the value of a stock with an expected annual dividend of $2.00 next year and estimated annual dividend growth of 2% per year indefinitely. Assume a discount rate of 8%. Solve the problem two different ways: first by using the algebraic formula for the Gordon Growth Model, then by using Excel to calculate and sum the dividends and their respective present values for the next 150 years. hint: Use the PV Const Growth Dividend example in the posted DDM Excel Examples file as a guide. Feel free to copy the worksheet and make the minor necessary changes to answer this question.
3. Calculate the value of a stock with the following expectations for dividend payments: $1.75 in Year 1, $2.00 in Year 2, and then annual dividend growth of 1.5% per year indefinitely. Assume a discount rate of 9%. Solve the problem two different ways: first by using the algebraic formula for the Gordon Growth Model combined with PV of uneven dividend payments, then by using Excel to calculate and sum the dividends and their respective present values for the next 150 years. hint: Use the Uneven, then Const. Growth Div example in the posted DDM Excel Examples file as a guide. Feel free to copy the worksheet and make the minor necessary changes to answer this question.
4. Calculate the value of a stock with the following expectations for dividend payments: $1.75 in Years 1, 2 and 3, and then annual dividend growth of 1.5% per year indefinitely. Assume a discount rate of 9%. Solve the problem two different ways: first by using the algebraic formula for the Gordon Growth Model combined with PV of uneven dividend payments, then by using Excel to calculate and sum the dividends and their respective present values for the next 150 years. hint: Use the Uneven, then Const. Growth Div example in the posted DDM Excel Examples file as a guide. Feel free to copy the worksheet and make the minor necessary changes to answer this question.
Answer #1

(1) Annual Preferred Stock Dividend = Dp = $ 2.45 and Discount Rate = kp = 9.5 %

Preferred Stock Price = Dp / kp = 2.45 / 0.095 = $ 25.789474 ~ $ 25.79

NOTE: Please raise separate queries for solutions to the remaining unrelated questions, as one query is restricted to the solution of only one complete question with up to four sub-parts.